OpenStudy (anonymous):

The linear density of a rod of length 4m is given by p(x)+5+4sqrt(x) measured in kilograms per meter, where x is measured in meters from end to end.find the mass.

jimthompson5910 (jim_thompson5910):

if m(x) represents the mass of the rod, where x is the length of the rod, then \[\Large \int p(x)dx = m(x)+C\]

jimthompson5910 (jim_thompson5910):

put another way \[\Large \frac{d}{dx}[m(x)] = \frac{dm}{dx} = p(x)\]

jimthompson5910 (jim_thompson5910):

hopefully that's enough to get you started